OpenStudy (anonymous):

Find the derivative of s(x) = -9 - 3x at x = 8

OpenStudy (anonymous):

The derivative of this is a constant, so it doesn't matter what x is.

OpenStudy (jhannybean):

\[\large s(x) =-3x-9\]\[\large s'(x)= \frac{d}{dx}(-3x-9)\]\[\large s'(x)=\frac{d}{dx}(-3x)-\frac{d}{dx}(9)\]\[\large s'(x)=-3\]

OpenStudy (anonymous):

ok thats what i was thinking thanx!!!!!!!!!

OpenStudy (jhannybean):

no problemo